ROTATIONAL SPECTRA OF ISOTOPIC CH_{3}CN IN THEIR v_{8} = 1 EXCITED VIBRATIONAL STATES

\underline{HOLGER S. P. MÜLLER}, I. Physikalisches Institut, Universität zu Köln, 50937 Köln, Germany; B. J. DROUIN, J. C. PEARSON, Jet Propulsion Laboratory, California Institute of Technology, Pasadena, CA 91109, USA; A. BELLOCHE, K. M. MENTEN, Max-Planck Institut für Radioastronomie, 53121 Bonn, Germany.

Methyl cyanide, CH_{3}CN, is an important interstellar molecule, in particular in hot and dense molecular cores, and it may play a role in the atmospheres of planets or of Titan. Therefore, we have recorded extensive rotational spectra up to ∼ 1.6 THz. Ground state rotational transitions of a number of minor isotopologs could be identified up to 1.2 THz in natural isotopic composition, including CH_{2}DCN and ^{13}CH_{3}^{13}CN.^{a} Recently, we have analyzed the rotational spectra of ^{13}CH_{3}CN, CH_{3}^{13}CN, and CH_{3}C^{15}N in their v_{8} = 1 excited vibrational states from spectra covering most of the frequencies between 0.44 THz and 1.20 THz. The analyses of the ^{15}N and ^{13}C species were facilitated by previous data up to 144 GHz and 56 GHz, respectively. Spectroscopic parameters determined in the fits will be compared with those of the main isotopolog.^{b} The importance of these results, in particular for radio-astronomical observations with the Atacama Large Millimeter Array (ALMA), will be stressed by the detection of transitions pertaining to the ^{13}C species in Sagittarius B2(N).

aH. S. P. Müller, B. J. Drouin, and J. C. Pearson, Astron. Astrophys. 506 (2009) 1187.

bH. S. P. Müller et al., manuscript in preparation.